# Sample-efficient and Scalable Exploration in Continuous-Time RL

### v1

@@ 1 Introduction
We propose COMBRL, a continuous-time model-based RL algorithm that uses probabilistic models, e.g., Gaussian processes and Bayesian neural networks, to capture the epistemic uncertainty of the learned model. Policies are selected by maximizing a weighted sum of extrinsic reward and epistemic uncertainty, following the optimism-in-the-face-of-uncertainty principle (Curi et al., 2020; Kakade et al., 2020; Treven et al., 2023; Sukhija et al., 2025b). In the unsupervised case, COMBRL actively learns the dynamics by guiding the agent toward regions where the model is most uncertain.
@@ 2 Problem setting
Optimizing solely for the reward function $r({\bm{x}}(t),{\bm{u}}(t))$ during learning introduces a directional bias, since $\bm{f}$$^{*}$ is mostly explored in high-reward regions of the state-action space, resulting in limited exploration and poor generalization of the learned dynamics. However, in many practical scenarios, such as unsupervised RL (Aubret et al., 2019; Sekar et al., 2020; Sukhija et al., 2023) or system identification (Åström and Eykhoff, 1971), the objective is not to optimize a performance criterion, but to learn the underlying nonlinear ODE $\bm{f}$$^{*}$ as fast and accurately as possible.
@@ 2.1 Related work
We briefly mention the most relevant related works below. An extended discussion of related works is provided in Appendix˜A.
Yildiz et al. (2021) introduce a continuous-time model-based RL approach that greedily maximizes the reward integral. Treven et al. (2023) show that greedy exploration performs suboptimally and propose leveraging optimistic dynamics to encourage exploration in uncertain regions. The resulting algorithm, OCORL, enjoys convergence guarantees under common continuity assumptions on the underlying ODE. However, the method does not scale to higher dimensional systems since it relies on an extremely challenging optimization over the set of plausible dynamics to enforce optimism.
Moreover, both aforementioned works rely on an external reward signal. When no extrinsic reward is available, or when accurate dynamics modeling is prioritized over extrinsic performance, these methods cannot be applied. On the other hand, intrinsic motivation techniques have long been used for this purpose (Pathak et al., 2017; Bellemare et al., 2016; Salge et al., 2014; Sekar et al., 2020). In particular, Sukhija et al. (2023) show convergence of intrinsic exploration methods in discrete time. However, their continuous-time counterpart is much less understood.
@@ 3.1 Optimistic planning objective
Earlier continuous-time optimistic methods addressed exploration by jointly optimizing over policies and dynamics ${\bm{f}}_{n}\in{\mathcal{M}}_{n-1}\cap{\mathcal{F}}$. This is intractable and heuristically addressed using a reparametrization trick from Curi et al. (2020), which increases input dimensionality from $d_{u}$ to $d_{u}+d_{x}$, limiting scalability in high-dimensional settings, e.g., control from pixels. COMBRL avoids this by selecting any model from ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$; in practice, using ${\bm{\mu}}_{n}$ works well. [^2] Even though the mean model might not lie in ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$. For GP dynamics, we show how to pick a model from ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$ in Section B.2. Unlike optimistic dynamics or classical control, COMBRL balances exploration and exploitation via a single scalar $\lambda_{n}$ and encourages exploration of uncertain regions to improve model fidelity. COMBRL is summarized in Algorithm˜1.
@@ 3.3 Theoretical results
Under regularity assumptions (Ass. 1 to 3), we have with probability at least $1-\delta$:
@@
Theorem˜1 bounds the regret of COMBRL w.r.t. the model complexity. For GP dynamics, where the well-calibration assumption is true and the monotonicity of the variance holds, Treven et al. (2023) show that the model complexity ${\mathcal{I}}_{N}({\bm{f}}^{*},S)$ is sublinear for many common kernels and MSSs, e.g., grows with $\operatorname{poly}\log(N)$ for the RBF kernel and equidistant MSS. Therefore, for common kernels and MSSs, COMBRL enjoys sublinear regret in the GP setting, and the policy converges to the optimal policy $\bm{\pi}$$^{*}$.
@@
Theorem˜2 provides a sample complexity bound for the unsupervised case. Effectively, this shows that pure intrinsic exploration with $\lambda_{n}\to\infty$ reduces our model epistemic uncertainty with a rate of $\sqrt{{{\mathcal{I}}^{3}_{N}}/{N}}$. To the best of our knowledge, we are the first to show this for continuous-time RL. We provide the proofs for all our theoretical results in Section˜B.2.
@@ 4 Experiments
We evaluate the COMBRL algorithm on various environments from the OpenAI/Farama gymnasium benchmark suite (Gym, Brockman et al. (2016); Towers et al. (2024)) and the DeepMind control suite (DMC, Tassa et al. (2018); Tunyasuvunakool et al. (2020)). For the dynamics model ${\bm{f}}_{n}$, we use Gaussian processes (GPs) and probabilistic ensembles (PEs) to capture uncertainty about well-calibrated statistical models. Since COMBRL models continuous-time dynamics directly, it also remains agnostic to the solver or discretization scheme, and can in principle accommodate adaptive strategies.
@@
Concretely, we first evaluate COMBRL with equidistant MSS and fixed control rates, solving the continuous-time planning problem using discrete-time solvers with fine-grained discretization to ensure accurate approximation of the underlying ODE. Unlike discrete-time RL, however, ${\bm{f}}_{n}$ learns the ODE of the true system $\bm{f}$$^{*}$. We then extend our evaluation to the time-adaptive setting using the sensing and control framework from Treven et al. (2024), comparing adaptive MSS against fixed-rate control.
For policy training and planning, we use SAC (Haarnoja et al., 2018) for closed-loop control and iCEM (Pinneri et al., 2021) for real-time trajectory optimization. For the time-adaptive control experiments, we use TaCoS (Treven et al., 2024) as an alternative planner. Additional details on the experimental setup are provided in Appendix˜C.
@@
We compare COMBRL with two baselines with different planning approaches. The mean planner uses the mean estimate ${\bm{\mu}}_{n}$ of the statistical model and greedily maximizes the extrinsic reward for the task at hand, akin to Yildiz et al. (2021). Furthermore, we compare our method with the trajectory sampling scheme (TS-1) proposed by Chua et al. (2018), subsequently referred to as PETS. PETS samples trajectories from PEs for state propagation and thus inherently captures the epistemic uncertainty during planning.
Figure 1: GP dynamics. Learning curves for baselines, COMBRL and OCORL with fixed internal reward weight $\lambda_{n}$ using GP dynamics and iCEM planning, averaged over 5 seeds. We report the mean and the standard error bands. COMBRL achieves higher asymptotic returns than PETS and mean, while matching or exceeding OCORL at roughly $3\times$ lower computational cost (Section˜C.2).
@@
Results, averaged over five random seeds with standard error bands, show that COMBRL consistently achieves higher asymptotic returns than the baselines, confirming the benefits of intrinsic rewards for guiding exploration. For the Pendulum experiments, all algorithms solve the task of swinging up. However, the co-optimization over the reward and the optimistic dynamics as well as the reparametrization trick used by the OCORL algorithm are computationally prohibitive, and require around $3\times$ the compute time compared to COMBRL (see Section˜C.2). Thus our algorithm scales better and is more computationally efficient than the state of the art.
Figure 2: Effect of intrinsic rewards. Learning curves for COMBRL (with auto-tuned $\lambda_{n}$) and baselines. We report the mean return when evaluating the learned model on the task at hand, averaged across 10 random seeds along with the standard error. COMBRL achieves the largest performance gains in sparse or underactuated tasks, and consistent improvements in higher-dimensional domains.
@@
To assess the effect of intrinsic rewards, we compare COMBRL with $\lambda_{n}>0$ against PETS and the mean planner ($\lambda_{n}=0$). Figure˜2 shows learning curves on several environments from Gym and DMC. In these experiments, we model dynamics using PEs and use SAC to train the policy. We compare COMBRL, with a nonzero intrinsic reward weight $\lambda_{n}$, against PETS and the mean planner. For COMBRL, the internal reward weight $\lambda_{n}$ is automatically tuned following the strategy proposed by Sukhija et al. (2025a), as outlined in Section˜C.3.
Across a range of environments, COMBRL achieves higher asymptotic returns, particularly in sparse-reward or underactuated settings such as MountainCar and CartPole, highlighting that optimism-driven exploration significantly accelerates learning. In higher-dimensional environments such as HalfCheetah and Hopper, COMBRL improves performance by encouraging exploration of uncertain regions, which helps uncover effective behaviors in complex, high-degree-of-freedom systems, even when these behaviors are not directly tied to high immediate reward signals.
@@
We evaluate whether exploration driven by model uncertainty improves generalization to unseen tasks. Specifically, we train policies on a primary task (e.g., reaching a target) and assess zero-shot performance on a downstream task not encountered during training (e.g., moving away from the target).
Figure˜3 compares standard COMBRL as well as its unsupervised variant ($\lambda_{n}\to\infty$, cf. Section˜3.2) to the baselines. To ensure a fair comparison, we let each agent explore the environment for several episodes, and then periodically evaluate the learned model on downstream tasks. While COMBRL generally achieves the best performance on the primary task especially in environments which offer sparse rewards, its unsupervised variant performs best on the downstream task across all seven evaluated domains. This suggests that exploration guided by model uncertainty encourages the agent to cover a more diverse state-action space and highlights the trade-off between task focus and exploration breadth: Ignoring the reward signal entirely leads the agent to explore broadly, acquiring a globally accurate model that generalizes better to unseen tasks for $\lambda_{n}\rightarrow\infty$.
@@
A key advantage of CT RL over its discrete-time counterpart is that it enables us to deploy policies at varying control frequencies. To illustrate this benefit, we extend our evaluation to the time-adaptive control and sensing (TaCoS) framework from Treven et al. (2024), where interactions with the system incur transition costs and the agent jointly optimizes both actions and their durations, i.e., the control frequency. We compare COMBRL-TaCoS with OTaCoS (the variant of OCORL for TaCoS), Mean-TaCoS, and PETS-TaCoS, as well as COMBRL with a fixed control rate and equidistant MSS. Across benchmarks, COMBRL-TaCoS achieves competitive or superior returns while using fewer interactions compared to the fixed control rate variant, which is shown by the higher returns in Figure˜5 and in turn, demonstrates sample efficiency. Compared to the previous experiments, this shows that optimism-driven exploration extends naturally to the adaptive-control setting and that COMBRL is, in principle, applicable independently of not only the dynamics model, but also the measurement selection strategy. We offer details on the TaCoS framework in Section˜C.6.
@@ 5 Conclusion
All algorithmic details of COMBRL, including the problem formulation, regret analysis, and assumptions, are presented in the main text, with complete proofs provided in Appendix˜B. The experimental setup, including environment details, model architectures, and hyperparameters, is described in Appendix˜C. The source code for the conducted experiments is publicly available on our project website: https://go.klem.nz/combrl, and on GitHub: https://github.com/lasgroup/ombrl.
@@ Acknowledgements
This project has received funding from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and the Microsoft Swiss Joint Research Center. Bhavya Sukhija was gratefully supported by ELSA (European Lighthouse on Secure and Safe AI) funded by the European Union under grant agreement No. 101070617. Numerical simulations were performed on the ETH Zürich Euler cluster.
@@ B.2 Analysis of Gaussian process dynamics
˜4 allows us to model $\bm{f}$$^{*}$ with GPs. The posterior mean ${\bm{\mu}}_{n}({\bm{z}})=[\mu_{n,j}({\bm{z}})]_{j\leq d_{\bm{x}}}$ and epistemic uncertainty ${\bm{\sigma}}_{n}({\bm{z}})=[\sigma_{n,j}({\bm{z}})]_{j\leq d_{\bm{x}}}$ can then be obtained using the following formula
@@
Consider the noise-free case, i.e., $\epsilon_{n,i}=0$, and let $\bar{{\bm{\mu}}}_{n}$ be the posterior mean for this setting. For the function $\bar{{\bm{\mu}}}_{n}$, it holds that $\left\lVert{\bm{f}}^{*}-\bar{{\bm{\mu}}}_{n}\right\rVert_{k_{n}}\leq\beta_{n}$ (Corollary 3.11 of Kanagawa et al. (2018)) and $\left\lVert\bar{{\bm{\mu}}}_{n}\right\rVert_{k}\leq B$ (Theorem 3.5 of Kanagawa et al. (2018)). Thus it follows that
@@
Hence the optimization problem Equation˜6 is equivalent to:
@@
Furthermore, let $\lambda_{n}=2\beta_{n}L_{r}(1+L_{{\bm{\pi}}})e^{L_{*}(1+L_{{\bm{\pi}}})T}$.
@@
$|J({\bm{\pi}},{\bm{f}}^{*})-J({\bm{\pi}},{\bm{f}}_{n})|$ $=\mathbb{E}\left[\int_{0}^{T}r({\bm{x}}(t),{\bm{\pi}}({\bm{x}}(t)))-r({\bm{x}}^{\prime}(t),{\bm{\pi}}({\bm{x}}^{\prime}(t)))\,dt.\right]$
$\leq L_{r}(1+L_{{\bm{\pi}}})\mathbb{E}\left[\int_{0}^{T}\left\lVert{\bm{x}}(t)-{\bm{x}}^{\prime}(t)\right\rVert\,dt.\right]$
$\leq 2\beta_{n}L_{r}(1+L_{{\bm{\pi}}})Te^{L_{\bm{f}}(1+L_{{\bm{\pi}}})T}\int_{0}^{T}\left\lVert{\bm{\sigma}}_{n-1}({\bm{x}}(t),{\bm{\pi}}({\bm{x}}(t)))\right\rVert\,dt.$ (Treven et al. (2023) Lemma 4.)
Similarly, we can also use Lemma 4 from Treven et al. (2023) and bound the regret with $\Sigma_{n}({\bm{\pi}},{\bm{f}}_{n})$.
$|J({\bm{\pi}},{\bm{f}}^{*})-J({\bm{\pi}},{\bm{f}}_{n})|$ $=\mathbb{E}\left[\int_{0}^{T}r({\bm{x}}(t),{\bm{\pi}}({\bm{x}}(t)))-r({\bm{x}}^{\prime}(t),{\bm{\pi}}({\bm{x}}^{\prime}(t)))\,dt.\right]$
$\leq L_{r}(1+L_{{\bm{\pi}}})\mathbb{E}\left[\int_{0}^{T}\left\lVert{\bm{x}}(t)-{\bm{x}}^{\prime}(t)\right\rVert\,dt.\right]$
$\leq 2\beta_{n}L_{r}(1+L_{{\bm{\pi}}})Te^{L_{B}(1+L_{{\bm{\pi}}})T}\int_{0}^{T}\left\lVert{\bm{\sigma}}_{n-1}({\bm{x}}^{\prime}(t),{\bm{\pi}}({\bm{x}}^{\prime}(t)))\right\rVert\,dt.$
@@
Let $\Sigma^{*}_{N}=\max_{{\bm{\pi}}}\Sigma_{N}({\bm{\pi}},{\bm{f}}^{*})$ and ${\bm{\pi}}^{*}_{N}$ the corresponding policy.
@@
$\leq\sum^{N}_{n=1}(1+\lambda_{n})^{2}\Sigma_{n}({\bm{\pi}}_{n},{\bm{f}}^{*})$ (Lemma 4)
@@
$\leq(1+\lambda_{N})^{2}\frac{1}{\sqrt{N}}\sum^{N}_{n=1}\Sigma^{2}_{n}({\bm{\pi}}_{n},{\bm{f}}^{*})$
$\leq{\mathcal{O}}\left(\sqrt{\frac{{\mathcal{I}}^{3}_{N}}{N}}\right)$

### v2

@@ 1 Introduction
We propose COMBRL, a continuous-time model-based RL algorithm that uses probabilistic models, e.g., Gaussian processes and Bayesian neural networks, to capture the epistemic uncertainty of the learned model. Policies are selected by maximizing a weighted sum of extrinsic reward and epistemic uncertainty, following the optimism-in-the-face-of-uncertainty principle (Curi et al., 2020; Kakade et al., 2020; Treven et al., 2023; Sukhija et al., 2025b). In the unsupervised setting often encountered in system identification (Åström and Eykhoff, 1971) and in unsupervised exploration (Aubret et al., 2019; Sekar et al., 2020; Sukhija et al., 2023), COMBRL reduces to active learning (Taylor et al., 2021), specifically uncertainty sampling (Lewis and Catlett, 1994) and guides the agent toward regions where the model is most uncertain.
@@
Unlike discrete-time RL, where interaction occurs on a fixed grid and uncertainty is evaluated only at visited steps, continuous-time reinforcement learning requires the agent to additionally decide when to observe and control the system. As a result, uncertainty accumulates along entire trajectories, and regret depends not only on the policy but also on the measurement schedule. Prior continuous-time approaches typically enforce optimism by optimizing jointly over policies and plausible dynamics, which is computationally demanding. In contrast, COMBRL realizes optimism through an intrinsic reward and directly trades off reward and epistemic uncertainty, yielding a scalable exploration mechanism whose guarantees explicitly depend on the measurement strategy.
@@ 2 Problem setting
Optimizing solely for the reward function $r({\bm{x}}(t),{\bm{u}}(t))$ during learning introduces a directional bias, since $\bm{f}$$^{*}$ is mostly explored in high-reward regions of the state-action space, resulting in limited exploration and poor generalization of the learned dynamics. However, in many practical scenarios, such as unsupervised RL or system identification, the objective is not to optimize a performance criterion, but to learn the underlying nonlinear ODE $\bm{f}$$^{*}$ as fast and accurately as possible.
@@ 2.1 Related work
We briefly mention the most relevant related works below. An extended discussion of related works is provided in Appendix A.
Yildiz et al. (2021) introduce a continuous-time model-based RL approach that greedily maximizes the reward integral. Treven et al. (2023) show that greedy exploration performs suboptimally and propose leveraging optimistic dynamics to encourage exploration in uncertain regions. The resulting algorithm, OCORL, enjoys convergence guarantees under common continuity assumptions on the underlying ODE. However, the method does not scale to higher dimensional systems since it relies on a challenging optimization over the set of plausible dynamics to enforce optimism.
Moreover, both aforementioned works rely on an external reward signal. When no extrinsic reward is available, or when accurate dynamics modeling is prioritized over extrinsic performance, these methods cannot be applied. On the other hand, intrinsic motivation techniques have long been used for this purpose (Salge et al., 2014; Bellemare et al., 2016; Pathak et al., 2017; Sekar et al., 2020). In particular, Sukhija et al. (2023) show convergence of intrinsic exploration methods in discrete time. However, their continuous-time counterpart is much less understood.
@@ 3.1 Optimistic planning objective
Earlier continuous-time optimistic methods addressed exploration by jointly optimizing over policies and dynamics ${\bm{f}}_{n}\in{\mathcal{M}}_{n-1}\cap{\mathcal{F}}$. This is intractable and heuristically addressed using a reparametrization trick from Curi et al. (2020), which increases input dimensionality from $d_{u}$ to $d_{u}+d_{x}$, limiting scalability in high-dimensional settings, e.g., control from pixels. COMBRL avoids this by selecting any model from ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$; in practice, using ${\bm{\mu}}_{n}$ works well. [^2] Even though the mean model might not lie in ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$. For GP dynamics, we show how to pick a model from ${\mathcal{M}}_{n-1}\cap{\mathcal{F}}$ in Appendix B.2. Unlike optimistic dynamics or classical control, COMBRL balances exploration and exploitation via a single scalar $\lambda_{n}$ and encourages exploration of uncertain regions to improve model fidelity. COMBRL is summarized in Algorithm˜1.
@@ 3.3 Theoretical results
Under regularity assumptions (Assumptions 1 to 3), the following holds with a probability of at least $1-\delta$:
@@
Theorem˜1 bounds the regret of COMBRL w.r.t. the model complexity. Importantly, this shows that the learning performance depends also on how efficiently the MSS collects informative measurements.
For GP dynamics, where the well-calibration assumption is true and the monotonicity of the variance holds, Treven et al. (2023) show that the model complexity ${\mathcal{I}}_{N}({\bm{f}}^{*},S)$ is sublinear for many common kernels and MSSs, e.g., grows with $\operatorname{poly}\log(N)$ for the RBF kernel and equidistant MSS. Therefore, for common kernels and MSSs, COMBRL enjoys sublinear regret in the GP setting, and the policy converges to the optimal policy $\bm{\pi}$$^{*}$.
@@
Theorem˜2 provides a sample complexity bound for the unsupervised case. Effectively, this shows that pure intrinsic exploration with $\lambda_{n}\to\infty$ reduces our model epistemic uncertainty with a rate of $\sqrt{{{\mathcal{I}}^{3}_{N}}/{N}}$. To the best of our knowledge, we are the first to show this for continuous-time RL.
We provide the proofs for all our theoretical results in Appendix B.
@@ 4 Experiments
We evaluate the COMBRL algorithm on various environments from the OpenAI/Farama gymnasium benchmark suite  (Gym, Brockman et al., 2016; Towers et al., 2024) and the DeepMind control suite (DMC, Tassa et al., 2018; Tunyasuvunakool et al., 2020). For the dynamics model ${\bm{f}}_{n}$, we use Gaussian processes (GPs) and probabilistic ensembles (PEs) to capture uncertainty about well-calibrated statistical models. Since COMBRL models continuous-time dynamics directly, it also remains agnostic to the solver or discretization scheme, and can in principle accommodate adaptive strategies.
@@
Concretely, we first evaluate COMBRL with equidistant MSS and fixed control rates, solving the continuous-time planning problem using discrete-time solvers with fine-grained discretization to ensure accurate approximation of the underlying ODE. Unlike discrete-time RL, however, ${\bm{f}}_{n}$ learns the ODE of the true system $\bm{f}$$^{*}$. We then extend our evaluation to the time-adaptive setting, comparing adaptive MSS against fixed-rate control.
We use ${\bm{f}}_{n}$ to generate simulated trajectories for policy training using soft actor-critic (SAC, Haarnoja et al., 2018) for closed-loop control and the improved cross-entropy method (iCEM, Pinneri et al., 2021) for real-time trajectory optimization. For the time-adaptive control experiments, we additionally employ the time-adaptive control & sensing (TaCoS) framework of Treven et al. (2024) as an alternative planner.
Brief descriptions of SAC and iCEM as well as additional implementation details are provided in Appendix C.1 and C.4, while further details on the TaCoS framework are given in Appendix C.7.
@@
We compare COMBRL with two baselines with different planning approaches. The mean planner uses the mean estimate ${\bm{\mu}}_{n}$ of the statistical model and greedily maximizes the extrinsic reward for the task at hand (i.e., $\lambda_{n}=0$), akin to Yildiz et al. (2021).
Furthermore, we compare our method with the trajectory sampling scheme (TS-1) proposed by Chua et al. (2018), subsequently referred to as PETS. PETS samples trajectories from PEs for state propagation and thus inherently captures the epistemic uncertainty during planning.
Figure 1: GP dynamics. Learning curves for baselines, COMBRL and OCORL with fixed internal reward weight $\lambda_{n}$ using GP dynamics and iCEM planning, averaged over 5 seeds. We report the mean and the standard error bands, and additionally the performance of MPC under the true dynamics as an estimate for the performance of the optimal policy. COMBRL achieves higher asymptotic returns than PETS and mean, while matching or exceeding OCORL at roughly $3\times$ lower computational cost (Appendix C.2).
@@
After each episode, we evaluate the learned model by computing the return using its mean prediction ${\bm{\mu}}_{n}$ on the original task. As a lower bound for the performance of the optimal policy $\bm{\pi}$$^{*}$, we report the best performance achieved by MPC (iCEM) on the true system when the dynamical system $\bm{f}$$^{*}$ is known. Thus, the MPC can plan with respect to the task/reward at hand subject to the true dynamics.
Results, averaged over five random seeds with standard error bands, show that COMBRL consistently achieves higher asymptotic returns than the baselines, confirming the benefits of intrinsic rewards for guiding exploration. For the Pendulum experiments, all algorithms solve the task of swinging up. However, the co-optimization over the reward and the optimistic dynamics as well as the reparametrization trick used by the OCORL algorithm are computationally prohibitive, and require around $3\times$ the compute time compared to COMBRL (see Appendix C.2). Thus our algorithm scales better and is more computationally efficient than the state of the art.
Figure 2: Effect of intrinsic rewards. Learning curves for COMBRL (with auto-tuned $\lambda_{n}$) and baselines on different continuous control tasks. We report the mean return when evaluating the learned model on the task at hand, averaged across 10 random seeds along with the standard error. COMBRL achieves the largest performance gains in sparse or underactuated tasks, and consistent improvements in higher-dimensional domains.
@@
To assess the effect of intrinsic rewards, we compare COMBRL with $\lambda_{n}>0$ against PETS and the mean planner. Figure˜2 shows learning curves on several environments from Gym and DMC. In these experiments, we model dynamics using PEs and use SAC to train the policy. We compare COMBRL with a nonzero intrinsic reward weight $\lambda_{n}$ against PETS and the mean planner ($\lambda_{n}=0$). For COMBRL, the internal reward weight $\lambda_{n}$ is automatically tuned following the strategy proposed by Sukhija et al. (2025a). Details of the tuning procedure and the resulting evolution of $\lambda_{n}$ over training are provided in Appendix C.3. We further show in Appendix C.6 that this auto-tuned intrinsic reward drives exploration towards epistemically uncertain regions of the state–action space.
Across a range of environments, COMBRL achieves higher asymptotic returns, particularly in sparse-reward or underactuated settings such as MountainCar and CartPole, highlighting that optimism-driven exploration significantly accelerates learning. In higher-dimensional environments such as HalfCheetah, Hopper, and Humanoid, COMBRL improves performance by encouraging exploration of uncertain regions, which helps uncover effective behaviors in complex, high-degree-of-freedom systems, even when these behaviors are not directly tied to high immediate reward signals.
@@
We evaluate whether exploration driven by model uncertainty improves generalization to unseen tasks. Specifically, we train policies on a primary task (e.g., reaching a target) and assess zero-shot performance on a downstream task not encountered during training (e.g., moving away from the target). The downstream tasks are implemented by modifying the reward per environment and are summarized in Appendix C.5.
Figure˜3 compares standard COMBRL as well as its unsupervised variant ($\lambda_{n}\to\infty$, cf. Section˜3.2) to the baselines. To ensure a fair comparison, we let each agent explore the environment for several episodes, and then periodically evaluate the learned model on downstream tasks. While COMBRL generally achieves the best performance on the primary task especially in environments which offer sparse rewards, its unsupervised variant performs best on the downstream task across all seven evaluated domains.
This suggests that exploration guided by model uncertainty encourages the agent to cover a more diverse state-action space and highlights the trade-off between task focus and exploration breadth: Ignoring the reward signal entirely leads the agent to explore broadly, acquiring a globally accurate model that generalizes better to unseen tasks for $\lambda_{n}\rightarrow\infty$.
@@
A key advantage of continuous-time RL over its discrete-time counterpart is that it enables us to deploy policies at varying control frequencies. To illustrate this benefit, we extend our evaluation to the time-adaptive control & sensing (TaCoS) framework from Treven et al. (2024), where interactions with the system incur transition costs and the agent jointly optimizes both actions and their durations, i.e., the control frequency. We offer details on the TaCoS framework in Appendix C.7.
We compare COMBRL with time-adaptive MSS (COMBRL-TaCoS) with OTaCoS (the variant of OCORL for TaCoS), Mean-TaCoS, and PETS-TaCoS. Furthermore, we compare these time-adaptive methods to COMBRL with a fixed control rate and equidistant MSS such as the one used in the previous experiments. Across benchmarks, COMBRL-TaCoS achieves competitive or superior returns while using fewer interactions compared to the fixed control rate variant, which is shown by the higher returns in Figure˜5 and in turn, demonstrates sample efficiency. Compared to the previous experiments, this shows that optimism-driven exploration extends naturally to the adaptive-control setting and that COMBRL is, in principle, applicable independently of not only the dynamics model, but also the measurement selection strategy.
@@ 5 Conclusion
LLM usage statement
Parts of this text were revised with the assistance of a large language model to aid or polish writing and to improve grammar and clarity; the authors remain responsible for all content.
@@
All algorithmic details of COMBRL, including the problem formulation, regret analysis, and assumptions, are presented in the main text, with complete proofs provided in Appendix B. The experimental setup, including environment details, model architectures, and hyperparameters, is described in Appendix C. The source code for the conducted experiments is available under https://github.com/lasgroup/ombrl.
@@ Acknowledgements
We would like to thank Bruce D. Lee for the insightful feedback on this work. This project has received funding from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. Numerical simulations were performed on the ETH Zürich Euler cluster.
@@ B.2 Analysis of Gaussian process dynamics
˜4 allows us to model $\bm{f}$$^{*}$ with GPs. The posterior mean ${\bm{\mu}}_{n}({\bm{z}})=[\mu_{n,j}({\bm{z}})]_{j\leq d_{\bm{x}}}$ and epistemic uncertainty ${\bm{\sigma}}_{n}({\bm{z}})=[\sigma_{n,j}({\bm{z}})]_{j\leq d_{\bm{x}}}$ can then be obtained using
@@
Consider the noise-free case, i.e., $\epsilon_{n,i}=0$, and let $\bar{{\bm{\mu}}}_{n}$ be the posterior mean for this setting. For the function $\bar{{\bm{\mu}}}_{n}$, it holds that $\left\lVert{\bm{f}}^{*}-\bar{{\bm{\mu}}}_{n}\right\rVert_{k_{n}}\leq\beta_{n}$ (Kanagawa et al., 2018, Corollary 3.11) and $\left\lVert\bar{{\bm{\mu}}}_{n}\right\rVert_{k}\leq B$ (Kanagawa et al., 2018, Theorem 3.5). Thus it follows that
@@
Hence the optimization problem in Equation˜6 is equivalent to:
@@
Furthermore, let $\lambda_{n}=2\beta_{n}L_{r}(1+L_{{\bm{\pi}}})Te^{L_{*}(1+L_{{\bm{\pi}}})T}$.
@@
We follow Treven et al. (2023) and bound the regret with $\Sigma_{n}({\bm{\pi}},{\bm{f}}_{n})$.
$|J({\bm{\pi}},{\bm{f}}^{*})-J({\bm{\pi}},{\bm{f}}_{n})|$ $=\mathbb{E}\left[\int_{0}^{T}r({\bm{x}}(t),{\bm{\pi}}({\bm{x}}(t)))-r({\bm{x}}^{\prime}(t),{\bm{\pi}}({\bm{x}}^{\prime}(t)))\,dt\right]$
$\leq L_{r}(1+L_{{\bm{\pi}}})\mathbb{E}\left[\int_{0}^{T}\left\lVert{\bm{x}}(t)-{\bm{x}}^{\prime}(t)\right\rVert\,dt\right]$
$\leq 2\beta_{n}L_{r}(1+L_{{\bm{\pi}}})Te^{L_{\bm{f}}(1+L_{{\bm{\pi}}})T}\int_{0}^{T}\left\lVert{\bm{\sigma}}_{n-1}({\bm{x}}(t),{\bm{\pi}}({\bm{x}}(t)))\right\rVert\,dt$ (Treven et al. (2023), Lemma 4)
@@
In the last line, we used the fact that $\Sigma_{n}(\cdot,\cdot)$ is bounded and positive, and thus behaves like a reward. In fact, it is an intrinsic reward (Sukhija et al., 2025b), which is why Lemma 4 also applies.
@@
Note that by Lemma 4, $\lambda_{N}$ scales with the confidence parameter $\beta_{N}$, which itself depends on the maximum information gain (Chowdhury and Gopalan, 2017; Rothfuss et al., 2023) and thus scales in the model complexity as $\lambda_{N}^{2}={\mathcal{O}}({\mathcal{I}}_{N})$, yielding the result. ∎
@@
Recall the definitions from Lemma 4. Let $\Sigma^{*}_{N}=\max_{{\bm{\pi}}}\Sigma_{N}({\bm{\pi}},{\bm{f}}^{*})$ and let ${\bm{\pi}}^{*}_{N}$ be the corresponding policy. Recall from Lemma 5 that $\Sigma_{n}(\cdot,\cdot)$ behaves like a reward.
@@
$\leq\frac{1}{N}\sum^{N}_{n=1}(1+\lambda_{n})^{2}\Sigma_{n}({\bm{\pi}}_{n},{\bm{f}}^{*})$ (Lemma 4)
@@
$\leq(1+\lambda_{N})^{2}\frac{\sqrt{N}}{N}\sqrt{\sum^{N}_{n=1}\Sigma^{2}_{n}({\bm{\pi}}_{n},{\bm{f}}^{*})}$ (Cauchy-Schwarz inequality)
$\leq{\mathcal{O}}\left(\sqrt{\frac{{\mathcal{I}}^{3}_{N}}{N}}\right)$ (c.f. Proof of Theorem 1)